Task

a) Evaluate each of the following expressions: 

i. 3(4−3)2 

ii. 3(5−3)2 

iii. 3(8−3)2

 iv. 3(10−3)2

b) In the expressions above, what parts are always the same? What part changes?

c) We can write an expression to represent all of the expressions in a) above if we use a letter for the part that changes. Write an expression using the letter x that could represent all of the numeric expressions in a).

d) What value of x would make your expression evaluate to 507? 


1
Expert's answer
2021-07-01T06:51:14-0400

a)

i.

"3(4-3)^2=3"

ii.

"3(5-3)^2=12"

iii.

"3(8-3)^2=75"

iv.

"3(10-3)^2=147"

b) In the expressions above the first multiplier "3" is the same, but the second multiplier changes from expression to expression.

In subtraction  the subtrahend "3" is the same, but the minuend changes from expression to expression.


c)


"3(x-3)^2"

d)


"3(x-3)^2=507"

"(x-3)^2=169"

"x-3=-\sqrt{169}\ \ or \ \ x-3=\sqrt{69}"

"x=3-13\ \ or \ \ x=3+13"

"x=-10\ \ or \ \ x=16"